Assume that ABC Company successfully issues a 6% convertible bond, due 12 years from now at $1,000 per bond. The bond pays interest 2 times per year. Also, assume that the bond is convertible into 100 shares of stock anytime during the life of the bond. Assume the common stock of the company is selling for $8 per share at the time the convertible bond is issued. The last assumption is that if ABC Company issued a regular bond (not convertible), they would have to sell the bond with a 9% coupon instead of the 6% on the convertible.

Using the information above, calculate the minimum value for the convertible if:

(A) ABC common stock trades at $14 per share

(B) The stock trades at $6 per share
Answer #1

a). If the share price is $14 then the bond will be trading 'off the stock'.

Its worth would be number of shares to be converted into*share price = 100*14 = $1,400

b). The share price falls to below the current price of $8.

Then, the bond has to be evaluated as an ordinary bond.

FV = 1,000; I = 9%/2 = 4.5% (coupon rate for an ordinary bond); N = 24; PMT = 30 (6%/2*1,000), solve for PV.

PV = 782.57

If the stock trades at $6 then the value of the bond will be $782.57